OpenStudy (anonymous):

Determine the slope of the secant line for the curve defined by the equation: f(x) = 3 x^2 + ( -5 ) for the points determined by x_0 = -6 and x_1 = -3. The slope of the secant line is m = .

OpenStudy (anonymous):

replace x by -6 and find \[f(-6)\] e then find \[f(-3)\] so you will then have two points \[(-6,f(-6))\] and \[(-3,f(-3))\] the slope will be \[\frac{f(-3)-f(-6)}{-3-(-6)}\] or \[\frac{f(-3)-f(-6)}{3}\]

OpenStudy (campbell_st):

f(-6) = 103 f(-3) = 22 points on the secant are (-6, 103) and (-3, 22) m = (103 -22)/(-6 --3) the gradient is m = -27

OpenStudy (anonymous):

thanks